octane burns completely in oxygen according to the equation
2C8H18 + 25O2 = 16CO2 =18H20
[a) calculate the volume of oxygen i.e. needed for the comp. combustion of 1 dm3 of octane vapor
b) what mass of water vapor will be formed
1
Expert's answer
2012-11-16T07:26:48-0500
V(C8H18): V(O2) = 2 : 25
V(O2) = 1 dm3 * 25/2 = 12.5 dm3
m = n * M;
n = V/ Vm
n(C8H18) = 1dm3/22.4 dm3/mol = 0.0446 mol;
n(H2O) = 0.0446mol *18/2 = 0.4 mol
m(H2O) = 0.4 mol * 18 g/mol = 7.22 g